# One-Step Diffusion-Based Image Compression with Semantic Distillation

### v1

@@ 4.1 Implementation and Evaluation Settings
Model and Training. Our generator adopts the U-Net architecture from the Stable Diffusion 1.5 (SD1.5) [44] and is initialized with weights from a pretrained one-step text-to-image task [58]. In Stage I, we use the pretrained tokenizer from MaskGIT [8] for semantic distillation. In Stage II, we employ the multi-step SD1.5 pretrained model as the teacher for diffusion distillation. To improve high resolution adaptability, we randomly crop patches of size 512 or 1024 during training. Models are optimized using AdamW [33]. Additional settings are provided in the supplementary material.
Evaluation datasets and metrics. We evaluate OneDC on several datasets, including Kodak [13], CLIC2020 test set [47], and MS-COCO 30K [30]. Reconstruction fidelity is assessed using reference perceptual metrics LPIPS [19] and DISTS [10], while generation realism is measured using the no-reference perceptual metric FID [16]. Bitrate saving is measured by BD-Rate [5]. It is worth noting that FID is not reported on the Kodak dataset due to its small size (24 images), which makes reliable FID computation infeasible [7]. More results are provided in the supplementary material.
Compared methods. We focus on comparing our OneDC with generative image codecs at ultra-low bitrate scenarios. The compared methods include the vanilla VAE-based generative codec MS-ILLM [39] and GLC [18], and the recent diffusion-based methods: PerCo [7], DiffEIC [28] and DiffC [49] (SD2.1 based). All methods are evaluated on the full resolution. Given that some diffusion-based codecs are evaluated on the CLIC2020 dataset with cropped $768\times 768$ resolution [7], we also provide the results on this setting in our supplementary material for better comparison.
@@ 4.2 Main Results
Quantitative Evaluation. Fig. 9 presents a quantitative comparison between OneDC and existing methods across multiple distortion metrics. Across all datasets—including high-resolution CLIC2020, low-resolution Kodak and MS-COCO 30K—OneDC consistently achieves the best reconstruction fidelity, as measured by LPIPS and DISTS. Compared with the previous SOTA multi-step diffusion-based DiffC, OneDC achieves a 50.08% bitrate saving on Kodak and 56.02% on CLIC2020 in terms of LPIPS quality, demonstrating its superior compression efficiency. In terms of generation realism, OneDC also achieves lower FID scores than all multi-step diffusion-based methods, ensuring visually pleasing, perceptually aligned reconstructions. On the MS-COCO 30K dataset, OneDC provides a 42.94% bitrate saving over PerCo—previously the highest-realism method—when evaluated by FID, further underscoring its effectiveness. In conclusion, OneDC consistently delivers faithful and perceptually compelling reconstructions across a variety of image resolutions and content complexities, despite relying on only a single sampling step. These results support our hypothesis that iterative sampling is not necessary for high-quality diffusion-based generative compression.
Qualitative Evaluation. Fig. 6 presents visual comparisons on the CLIC2020 test set. At the lowest bitrate, OneDC produces the most natural reconstructions with sharp high-frequency details, while MS-ILLM yields noticeably blurred textures. Multi-step diffusion-based methods (DiffEIC, PerCo, and DiffC) introduce structural distortions and deviated details. Similar trends are observed on the MS-COCO 30K dataset (Fig. 7), where MS-ILLM remains blurry and multi-step diffusion methods struggle to preserve visual fidelity. Additional examples are provided in the supplementary material.
[…]
@@
[…]
@@ 4.3 Ablation Studies and Efficiency Evaluation
Hybrid-domain Perceptual Learning. We evaluate the effectiveness of hybrid pixel-latent domain training in Stage II, as summarized in Table 1. (1) Using the pixel-domain loss only (No $L_{\text{distill}}$ and $\gamma L_{adv}$) leads to notable drops in both FID and DISTS scores. Visual examples in Fig. 10-Left show grid-like artifacts, reflecting a significant decline in perceptual quality. This underscores the importance of diffusion distillation in enforcing distributional alignment for realistic reconstruction. (2) Using the latent-domain loss only (No $L_{\text{recon}}$) also results in consistent degradation across all metrics, accompanied by visible color shifts in the reconstructions (Fig. 10-Right). This highlights the importance of pixel-level supervision in image compression, which emphasizes accurate reconstruction—unlike the pure image generation task that focuses solely on visual realism.
Operation Efficiency. We evaluate the coding times of different methods on $1024\times 1024$ images using an A100 GPU. These runtime results are presented alongside BD-Rate results on the MS-COCO 30K dataset to offer a comprehensive view of operational efficiency. As shown in Table 2, OneDC achieves a significant decoding speed advantage—over 20×20$\times$20 × faster—compared to multi-step diffusion-based methods, while still maintaining impressive reconstruction quality. Although OneDC is slightly slower than the pure VAE-based MS-ILLM due to its larger model, this added capacity is crucial for preserving perceptual quality under extreme compression, as evidenced by its superior DISTS and FID performance. Overall, these results underscore the effectiveness of OneDC in balancing fidelity, realism, and efficiency at ultra-low bitrates.
@@ 5 Conclusion
We present OneDC, a novel one-step diffusion-based generative image codec designed for high-quality and efficient compression under ultra-low bitrate scenarios. OneDC integrates a latent compression module for efficient coding and a one-step diffusion generator for fast reconstruction. To provide the critical semantic guidance for the one-step diffusion model and address the limitations of textual guidance, we leverage the hyperprior as the substitute. Furthermore, we improve the semantic capability of the hyperprior through a distillation mechanism. To enhance training efficiency and stability, we adopt a two-stage strategy: Stage I focuses on compression training through pixel-domain supervision, while the Stage II refines reconstruction quality via hybrid-domain perceptual learning. Extensive experiments demonstrate that OneDC achieves SOTA perceptual quality with fast decoding, highlighting the strong potential of one-step diffusion models in generative image compression.
@@ Appendix A Experiment
Evaluation of third-party models. We evaluate MS-ILLM [39] using the official checkpoints and fine-tune them with the provided code to support lower bitrates. For GLC [18], we report the results directly from its paper because the code and models are not publicly available at present. For DiffEIC [28] and DiffC [49], we use the official implementations and released models. Specifically, we adopt the Stable Diffusion (SD) 2.1-based variant of DiffC, which shows slightly better performance compared to the SD 1.5 version. For PerCo [7], we use a publicly available reimplementation [23], as the original code has not been released. For multi-step diffusion codecs, we follow the default sampling settings provided in their code (e.g., 50 for DiffEIC, 20 for PerCo).
Test setting of FID metric. For the CLIC2020 test set (both full resolution and the $768\times 768$ cropped variant), as well as the DIV2K test set [1], we compute FID using $256\times 256$ patches, following the protocol in [36]. This setting is also consistent with that used in DiffEIC [28]. For the MS-COCO 30K dataset, we use $512\times 512$ patches for FID evaluation, following the setup in PerCo [7].
Additional datasets and metrics. To enable a more comprehensive comparison, we additionally evaluate our method on the DIV2K test set [1]. Notably, several diffusion-based codecs [28, 49] report results on the CLIC2020 test set [47] at a resolution of $768\times 768$ [28]. To ensure fair comparison, we include evaluations under this setting as well. Performance on DIV2K and CLIC2020 ($768\times 768$) is presented in Fig. 9, where OneDC consistently achieves SOTA results across datasets and metrics, demonstrating strong robustness and generalization.
We also report traditional pixel-level distortion metrics—PSNR and MS-SSIM [52]—to provide a more comprehensive analysis, as shown in Fig. 10. At extremely low bitrates, optimizing for PSNR often suppresses high-frequency details, resulting in blurred reconstructions [36]. While OneDC shows slightly lower PSNR than MS-ILLM, both perceptual metrics and qualitative examples clearly demonstrate its superior visual quality. Despite prioritizing perceptual realism, OneDC still achieves competitive PSNR compared to other diffusion-based methods. Moreover, on the MS-SSIM metric, OneDC matches MS-ILLM and outperforms all other baselines. These results confirm that OneDC delivers strong pixel-level fidelity alongside high perceptual realism, highlighting the overall effectiveness of our method.
@@
Figure 9: Comparison of methods on the CLIC2020 test ($768\times 768$) and the DIV2K dataset.
Figure 10: Comparison of methods measured by PSNR and MS-SSIM.
@@ Appendix B Training Details
$L_{stageI}=L_{recon}+\lambda R+\alpha L_{aux},\;\;\;\;\text{where}\;\;L_{recon}=L_{1}+L_{perceptual}$ (5)
We use the L1 as the pixel-level loss and the LPIPS [19] as the perceptual-level loss. To support various bitrates, the rate-distortion trade-off parameter $\lambda$ is set to $\{1.8,2.9,4.6,7.4\}$. An auxiliary code prediction loss $L_{aux}$ is included with a weighting factor of $\alpha=0.001$. We train our model on the dataset introduced in [14]. Training is performed on 4×A100 GPUs for 800,000 steps, using a three-stage learning rate schedule with AdamW [33]: a) 5e-5 for the first 500,000 steps; b) 1e-5 for the next 200,000 steps; c) 1e-6 for the final 100,000 steps. During training, image patches of size $\{512,1024\}$ are randomly cropped with probabilities of $\{0.6,0.4\}$, respectively. The batch size is set to 32 for $512\times 512$ crops and 8 for $1024\times 1024$ crops (across 4 GPUs).
@@
$L_{stageII}=L_{distill}+\beta L_{recon}+\gamma L_{adv},\;\;\;\;\text{where:}$: (6)
@@
$L_{fake}=\mathbb{E}_{t,\tilde{y}^{\prime}_{t}}||\epsilon_{fake}(\tilde{y}^{\prime}_{t},t)-\tilde{y}_{0}||^{2}_{2}$ (11)
@@
Training is conducted on 4×A100 GPUs for 1,000,000 steps. The learning rate is fixed at 1e-6 (with AdamW) for the one-step generator, fake network, and discriminator. The batch size, cropping strategy and training data are identical to those used in Stage I.
@@ Appendix C Model Details
For parameter-efficient adaptation, we insert LoRA [17] layers across all modules of the pretrained one-step diffusion U-Net, setting the LoRA rank to 64. The combined parameter count of the encoder and decoder components ($g_{a},h_{enc},h_{ctx},h_{sem}$ and $g_{s}$) is 394M, while the adapted diffusion U-Net contributes an additional 928M parameters.

### v2

@@ 4.1 Implementation and Evaluation Settings
Model and Training. Our generator adopts the U-Net architecture from the Stable Diffusion 1.5 (SD1.5) [50] and is initialized with weights from a pretrained one-step text-to-image task [66]. In Stage I, we use the pretrained tokenizer from MaskGIT [10] for semantic distillation. In Stage II, we employ the multi-step SD1.5 pretrained model as the teacher for diffusion distillation. To improve high-resolution adaptability, we randomly crop patches of size 512 or 1024 during training. Models are optimized using AdamW [38]. Additional settings are provided in the supplementary material.
Compared methods. We compare our OneDC with generative image codecs at a low-bitrate scenario. The compared methods include the vanilla VAE-based generative codec MS-ILLM [44] and GLC [22], and the recent diffusion-based methods: PerCo (SD) [28], DiffEIC [33], DiffC [57] (SD2.1-based) and DDCM [46]. For most methods, we evaluate at full-resolution, consistent with previous implementations in generative neural codecs [41, 44, 22]. For some diffusion-based methods like DDCM, we additionally provide results under the resize & center-crop setting, following DDCM protocol [46] for fair comparison. Further details are provided in the supplementary material.
Evaluation datasets and metrics. We evaluate OneDC on several datasets, including Kodak [16], CLIC2020 test set [55], and MS-COCO 30K [35]. Reconstruction fidelity is evaluated using perceptual metrics LPIPS [24] and DISTS [13], along with the traditional metrics PSNR and MS-SSIM [60], while generative realism is measured by the no-reference perceptual metric FID [19]. Bitrate saving is measured by BD-Rate [6]. It is worth noting that FID is computed on overlapping $256\times 256$ patches at full-resolution CLIC2020, following previous practice [44, 22]. For the MS-COCO 30K dataset with $512\times 512$ images, we evaluate FID on full-images, consistent with [9, 28]. However, large patch leads to unreliable FID on the small test set (Kodak with 24 images) [9]. To address this, under the resize & center-crop setting we follow the DDCM  [46] and compute FID on overlapping $64\times 64$ patches for Kodak. More details are provided in the supplementary material.
@@ 4.2 Main Results
Quantitative Evaluation. Fig. 7 presents the full-resolution comparison between OneDC and existing methods across multiple distortion metrics. Across all datasets, including high-resolution CLIC2020, low-resolution Kodak and MS-COCO 30K, OneDC consistently achieves the best reference-based perceptual fidelity, as measured by LPIPS and DISTS. Compared with the previous SOTA, multi-step diffusion-based DiffC, OneDC achieves 55.27% bitrate saving on Kodak and 54.60% on CLIC2020 in terms of LPIPS, demonstrating superior compression efficiency. On traditional metrics such as PSNR, OneDC also exhibits competitive performance. For generation realism, OneDC attains lower FID than all multi-step diffusion baselines when FID is evaluated on the CLIC2020 test set. On the MS-COCO 30K dataset, OneDC provides a 39.55% bitrate saving over PerCo (SD), previously the highest-realism method on this dataset. Such result further underscoring our method’s effectiveness. Furthermore, Fig. 8 reports results under the resize & center-crop setting: OneDC achieves better LPIPS and PSNR than DDCM while maintaining competitive realism, demonstrating strong fidelity-oriented perceptual compression. In conclusion, OneDC delivers faithful and perceptually compelling reconstructions across a wide range of resolutions and content complexities, despite relying on a single sampling step. These evaluation results support our hypothesis that iterative sampling is unnecessary for high-quality diffusion-based generative compression.
Qualitative Evaluation. Fig. 5 presents visual comparisons on the CLIC2020 test set (full-resolution). At the lowest bitrate, OneDC produces the most natural and faithful reconstructions with sharp high-frequency details, while MS-ILLM yields noticeably blurred textures and multi-step diffusion-based methods (DiffEIC, PerCo, and DiffC) introduce structural distortions and inconsistent details. Similar trends are observed on the MS-COCO 30K dataset (Fig. 6), where MS-ILLM remains blurry and multi-step diffusion methods struggle to preserve visual fidelity. These qualitative results, together with the quantitative comparisons, highlight the SOTA performance of the proposed OneDC. Additional visual examples are provided in the supplementary material.
Table 1: Ablation studies with BD-Rate ($\%$) $\downarrow$.
@@
Table 2: Comparison of coding time and BD-Rate ($\%$) $\downarrow$.
@@
Notes: Metrics in the tables denote those used for BD-Rate calculation [6]; Ours is the anchor (0.00%).
@@ 4.3 Ablation Studies and Efficiency Evaluation
Hybrid-domain Perceptual Learning. We evaluate the effectiveness of hybrid pixel-latent domain training in Stage II, as summarized in Table 1. (1) Using the pixel-domain loss only (No $L_{distill}$ and $L_{adv}$) leads to notable drops in both FID and DISTS scores. Visual examples in Fig. 9-Left show grid-like artifacts, reflecting a significant decline in perceptual quality. This underscores the importance of diffusion distillation in enforcing distributional alignment for realistic reconstruction. (2) Using the latent-domain loss only (No $L_{recon}$) also results in consistent degradation across all metrics, accompanied by visible color shifts in the reconstructions (Fig. 9-Right). This highlights the importance of pixel-level supervision in image compression, which emphasizes accurate reconstruction—unlike the pure image generation task that focuses solely on visual realism.
Operation Efficiency. We evaluate the coding times of different methods on $1024\times 1024$ images using an A100 GPU. These runtime results are presented alongside BD-Rate results on the MS-COCO 30K dataset to offer a comprehensive view of operational efficiency. As shown in Table 2, OneDC achieves a significant decoding speed advantage—over $20\times$ faster—compared to multi-step diffusion-based methods, while still maintaining impressive reconstruction quality. Although OneDC is slightly slower than the pure VAE-based MS-ILLM due to its larger model, this added capacity is crucial for preserving perceptual quality under extreme compression, as evidenced by its superior DISTS and FID performance. Overall, these results underscore the effectiveness of OneDC in balancing fidelity, realism, and efficiency at ultra-low bitrates.
@@ 5 Conclusion
We present OneDC, a novel one-step diffusion-based generative image codec designed for high-quality and efficient compression under ultra-low bitrate scenarios. OneDC integrates a latent compression module for efficient coding and a one-step diffusion generator for fast reconstruction. To provide the critical semantic guidance for the one-step diffusion model and address the limitations of textual guidance, we leverage the hyperprior as a substitute. Furthermore, we improve the semantic capability of the hyperprior through a distillation mechanism. To enhance training efficiency and stability, we adopt a two-stage strategy: Stage I focuses on compression training through pixel-domain supervision, while Stage II refines reconstruction quality via hybrid-domain perceptual learning. Extensive experiments demonstrate that OneDC achieves SOTA perceptual quality with fast decoding, highlighting the strong potential of one-step diffusion models in generative image compression.
@@ Appendix A Experiment
Evaluation of third-party models. We evaluate MS-ILLM [44] using the official checkpoints and fine-tune them with the provided code to support lower bitrates. For GLC [22], we report the results directly from its paper because the code and models are not publicly available at present. For DiffEIC [33] and DiffC [57], we use the official implementations and released models. Specifically, we adopt the Stable Diffusion (SD) 2.1-based variant of DiffC, which shows slightly better performance compared to the SD 1.5 version. For PerCo [9], we use a publicly available reimplementation PerCo (SD) [28], as the original code has not been released. We also report the comparison with original PerCo by extract data from their paper (Fig. 10). For DDCM [46], we use the number reported in their paper. For multi-step diffusion codecs, we follow the default sampling settings provided in their code (e.g., 50 for DiffEIC, 20 for PerCo).
Test settings and FID calculation. At the full-resolution setting, we compute FID using overlapping $256\times 256$ patches for the CLIC2020 and DIV2K test sets [1], following the protocol of [41]. For the MS-COCO 30K dataset with $512\times 512$ images, FID is evaluated on entire images, consistent with [9, 28]. At the resize & center-crop setting, we resize the short side of each image (512 for Kodak, 768 for CLIC2020 test set) and then apply a center crop. In this setting, we use $64\times 64$ patches for FID calculation on Kodak and $128\times 128$ patches on CLIC2020 test set, consistent with DDCM [46]. It is worth noting that DiffEIC also evaluates under the resize & center-crop setting on CLIC2020 test set at 768 resolution, but computes FID with overlapping $256\times 256$ patches. For completeness, we additionally report results under DiffEIC’s protocol, denoted as resize & center-crop, 256 FID.
Additional datasets and metrics. To enable comprehensive comparison, we further evaluate our method on the DIV2K test set [1] under the full-resolution setting, as shown in Fig. 11. We also report results on the CLIC2020 test set at 768 resolution under DiffEIC’s protocol, i.e., resize & center-crop, 256 FID, as illustrated in Fig. 12. Across datasets, metrics, and evaluation settings, the OneDC consistently achieves SOTA performance, demonstrating strong robustness and generalization. The raw data used for evaluation is available at Table. 5 $\sim$  10.
We also report traditional pixel-level distortion metrics (PSNR and MS-SSIM [60]) to provide a more comprehensive analysis, as shown in Fig. 13. At extremely low bitrates, optimizing for PSNR often suppresses high-frequency details, resulting in blurred reconstructions [41]. While OneDC shows slightly lower PSNR than MS-ILLM, both perceptual metrics and qualitative examples clearly demonstrate its superior visual quality. Despite prioritizing perceptual quality, OneDC still achieves competitive PSNR compared to other diffusion-based methods. Moreover, on the MS-SSIM metric, OneDC matches MS-ILLM and outperforms all other baselines. These results confirm that OneDC delivers strong pixel-level fidelity alongside high perceptual realism, highlighting the overall effectiveness of our method.
@@
Bitrate allocation between hyperprior $\hat{z}$ and latent $\hat{z}$. We ablate the roles of hyperprior $\hat{z}$ and latent $\hat{y}$ on CLIC2020 test set by fixing the hyperprior budget to 0.0035 bpp and gradually increasing the bitrate of $\hat{y}$. Table 4 reports the results and reveals a clear division of each part’s role: (a) with zero bits allocated for latent $\hat{y}$, the codec still produces coherent reconstructions (also show in Fig. 20), indicating that $\hat{z}$ supplies a strong semantic information; (b) as bits are assigned to $\hat{y}$, fidelity and perceptual quality improve monotonically (e.g., LPIPS drops from 0.290 to 0.119), confirming that $\hat{y}$ is essential for encoding fine-grained details.
Table 3: Complexity analysis with model size. BD-Rate is calculated on the MS-COCO 30K dataset.
Model Params Enc. Time (s) Dec. Time (s) BD-Rate (Metircs)↓
LPIPS DISTS FID
MS-ILLM 181M 0.14 0.17 138.3% 253.0% 478.4%
DiffEIC 1.4B 0.32 12.4 305.0% 239.1% 341.0%
PerCo (SD) 3.8B+340M+955M* 0.58 8.80 538.8% 345.8% 59.6%
[…]
DDCM 950M - - - - -
OneDC 1.4B 0.15 0.34 0.00% 0.00% 0.00%
* Open-sourced PerCo includes an additional 3.8B BLIP2 caption model and 340M CLIP model.
Table 4: Bitrate allocation analysis on the CLIC2020 dataset.
Bpp $\hat{z}$ (ratio) Bpp $\hat{y}$ (ratio) Bpp Total PSNR↑ MS-SSIM↑ LPIPS↓ DISTS↓ FID↓
0.0035 (100%) 0.0 (0%) 0.0035 19.31 0.629 0.290 0.169 14.885
0.0035 (43%) 0.0047 (57%) 0.0082 23.13 0.790 0.163 0.089 6.223
0.0035 (27%) 0.0094 (73%) 0.0129 24.20 0.826 0.139 0.077 5.560
0.0035 (18%) 0.0157 (82%) 0.0192 25.25 0.856 0.119 0.068 4.979
Model size. We provide parameter counts, runtime comparisons (on 1024×1024 images), and BD-Rate results on the MS-COCO 30K dataset for better efficiency evaluation, as shown in Table 3. Compared to MS-ILLM, diffusion-based methods typically use larger models but achieve superior perceptual quality (e.g., lower BD-Rate with FID) due to stronger generative capacity. Unlike other diffusion-based codecs, OneDC avoids external caption models and multi-step sampling, enabling over 20× faster decoding while also achieving better rate-distortion performance.
Memroy usage. We also report memory usage under the resize & center-crop setting on $512\times 512$ Kodak images. PerCo (SD) requires about 22,220 MB of GPU memory, whereas DDCM uses 4,186 MB and OneDC uses 8,038 MB. These results highlight that incorporating a large language model, as in PerCo, substantially increases computational burden. DDCM achieves lower memory consumption by employing the SD model in a zero-shot manner. Although OneDC requires more memory than DDCM, its single-step design reduces inference cost compared to both multi-step PerCo and DDCM.
Figure 10: Comparison with original PerCo [9] on the MS-COCO 30K dataset.
Figure 11: Rate–distortion curves on the DIV2K dataset at full-resolution setting.
Figure 12: Rate–distortion curves on the CLIC2020 at resize & center-crop, 256 FID setting, with resized resolution 768.
Figure 13: Comparison of methods measured by PSNR and MS-SSIM at full-resolution setting.
@@ Appendix B Training Details
[…]
We use the L1 as the pixel-level loss and the LPIPS [24] as the perceptual-level loss. To support various bitrates, the rate-distortion trade-off parameter $\lambda$ is set to $\{0.6,1.0,1.8,2.9,4.6,7.4,12.2\}$. An auxiliary code prediction loss $L_{aux}$ is included with a weighting factor of $\alpha=0.001$. We train our model on the dataset introduced in [17]. Training is performed on 4×A100 GPUs for 800,000 steps, using a three-stage learning rate schedule with AdamW [38]: a) 5e-5 for the first 500,000 steps; b) 1e-5 for the next 200,000 steps; c) 1e-6 for the final 100,000 steps. During training, image patches of size $\{512,1024\}$ are randomly cropped with probabilities of $\{0.6,0.4\}$, respectively. The batch size is set to 32 for $512\times 512$ crops and 8 for $1024\times 1024$ crops (across 4 GPUs). This stage takes approximately 6 days as we use high-resolution patches for training.
@@
$L_{stageII}=L_{distill}+\beta L_{recon}+\gamma L_{adv},\;\;\;\;\text{where:}$ (7)
@@
[…]
@@
Training is conducted on 4×A100 GPUs for 1,000,000 steps. This stage requires around 12 days, as diffusion distillation introduces additional cost in addition to high-resolution training. The learning rate is fixed at 1e-6 (with AdamW) for the one-step generator, fake network, and discriminator. The batch size, cropping strategy and training data are identical to those used in Stage I.
@@ Appendix C Model Details
For parameter-efficient adaptation, we insert LoRA [21] layers across all modules of the pretrained one-step diffusion U-Net, setting the LoRA rank to 64. The combined parameter count of the encoder and decoder components ($g_{a},h_{enc},h_{ctx},h_{sem}$ and $g_{s}$) is 394M, while the adapted diffusion U-Net contributes an additional 928M parameters (about 860M for SD1.5 and 68M for LoRA).